# derivation

### v1

@@ A.1 Ablation Study on Datasets
To determine the optimal dataset for training binarized models, we conducted a comparative analysis using various training datasets, as summarized in Table 5. The results indicate that models trained solely on the WikiText2 dataset, due to its relatively small dataset size, tend to exhibit overfitting tendencies and struggle to generalize to other datasets. While these models demonstrate considerable perplexity improvement on the WikiText2 evaluation, their perplexity on the C4 dataset and zero-shot accuracy is notably poor. Conversely, models trained exclusively on the C4 dataset perform well across a wide range of tasks, except for the evaluation on WikiText2. Following the approach of previous research [16], we also experiment with a generated dataset synthesized using the LLaMA-1-7B model. Although this dataset generally performs satisfactorily across various language modeling tasks, its performance lags behind that of the C4 dataset. Therefore, to enhance overall model performance, we opt to train the models on a mixed dataset comprising both C4 and WikiText2. Moreover, the accessibility of both C4 and WikiText2 as open-source datasets further facilitates their adoption for training purposes.
@@
Datasets Perplexity $\downarrow$ Zero-shot Accuracy $\uparrow$
Wiki2 C4 BoolQ PIQA HellaSwag WinoGrande ARC-e ARC-c Avg

### v2

@@ 6 Conclusion
Acknowledgements
This work was supported in part by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. 2021-0-00105: Development of model compression framework for scalable on-device AI computing on Edge applications , IITP-2023-RS-2023-00256081: artificial intelligence semiconductor support program to nurture the best talents, No. 2021-0-01343: Artificial Intelligence Graduate School Program (Seoul National University) , No.2021-0-02068: Artificial Intelligence Innovation Hub), in part by the Samsung Research Funding Center under Project SRFC-TC1603-53, and BK21 FOUR program at Seoul National University.
@@ A.1 Ablation Study on Datasets
To determine the optimal dataset for training binarized models, we conduct a comparative analysis using various training datasets, as summarized in Table 5. The results indicate that models trained solely on the WikiText2 dataset, due to its relatively small dataset size, tend to exhibit overfitting tendencies and struggle to generalize to other datasets. While these models demonstrate considerable perplexity improvement on the WikiText2 evaluation, their perplexity on the C4 dataset and zero-shot accuracy is notably poor. Conversely, models trained exclusively on the C4 dataset perform well across a wide range of tasks, except for the evaluation on WikiText2. Following the approach of previous research [16], we also experiment with a generated dataset synthesized using the LLaMA-1-7B model. Although this dataset generally performs satisfactorily across various language modeling tasks, its performance lags behind that of the C4 dataset. Therefore, to enhance overall model performance, we opt to train the models on a mixed dataset comprising both C4 and WikiText2. Moreover, the accessibility of both C4 and WikiText2 as open-source datasets further facilitates their adoption for training purposes.
@@
Training Dataset Perplexity $\downarrow$ Zero-shot Accuracy $\uparrow$
Wiki2 C4 BoolQ PIQA Hella. WinoG. ARC-e ARC-c Avg
@@
A.2 Latency Measurement
To assess the latency of GEMV operation for our BinaryMoS, we have evaluated the latency of previous binarized models and the BinaryMoS by developing appropriate CUDA kernels for 1-bit matrix multiplication, modifying the CUDA kernel for multi-bit matrix multiplication [39]. Additionally, we further customize the CUDA kernel of BinaryMoS to fuse scaling experts and routing operations on top of the 1-bit matrix multiplication CUDA kernel. We measure the latency of the linear layers in LLaMA-7B and LLaMA-13B with batch size of 1 and results are presented in Table 6. All experiments are conducted on NVIDIA A6000 GPUs.
Previous methods like PB-LLM and BiLLM require extra matrix multiplications, making them very slow. OneBit, which employs the simplest binarization scheme, achieves significant improvement over the original Float16 model and shows the minimum latency. Meanwhile, our BinaryMoS introduces additional operations for processing scaling experts, which require far fewer operations compared to matrix multiplication. Consequently, BinaryMoS also shows similar latency results to OneBit. This demonstrates that the multi-scaling factor module in BinaryMoS improves performance in terms of perplexity and zero-shot accuracy with minimal overhead to latency.
Table 6: Latency ($\mu$sec) of linear layer in LLaMA-1/2-7B and LLaMA-1/2-13B.
Model Config LLaMA-1/2-7B LLaMA-1/2-13B
Weight Size 4096 $\times$ 4096 4096 $\times$ 11008 11008 $\times$ 4096 5120 $\times$ 5120 5120 $\times$ 13824 13824 $\times$ 5120
Float16 68.2 151.7 143.5 95.6 224.1 213.6
PB-LLM 96.1 177.5 168.3 122.7 243.7        234.7
BiLLM 87.1 96.4 104.2 95.2 124.2 131.0
OneBit 32.7 33.7 34.9 33.4 41.4 42.6
BinaryMoS 34.5 36.9 37.0 35.6 43.4 44.5
A.3 Experimental Results for LLaMA-1-30B
Table 7 provides further experimental results on the LLaMA-1-30B model. In line with the trends observed in Table 3, BinaryMoS consistently surpasses other binarization approaches for this 30B model. This assessment highlights the effectiveness of BinaryMoS for large-scale LLMs.
Table 7: Perplexity and zero-shot accuracy results of Float16 and binarized LLMs for LLaMA-1-30B
Wiki2 C4 BoolQ PIQA Hella. WinoG. ARC-e ARC-c Average
LLaMA-1-30B Float16 16 4.10 5.98 68.37 80.08 79.22 72.69 58.96 45.39 67.45
PB-LLM 1 32.24 34.59 62.18 60.22 33.75 53.91 33.50 24.40 44.66
BiLLM 1 10.10 12.28 62.39 71.38 59.49 66.54 44.95 33.19 56.32
BinaryMoS 1 6.63 8.36 67.76 75.62 67.38 64.09 49.62 35.07 59.92
A.4 Generation Quality
Figure 4 compares the generation quality between BinaryMoS and OneBit on LLaMA-1-13B model. BinaryMoS can generate contextually proper answers, whereas OneBit fails to generate correct answers. These results demonstrate that the BinaryMoS processes each token with token-adaptive scaling factors which contain contextual information and the improvement of BinaryMoS over OneBit is substantial enough to make binarized LLMs more applicable in practice.
Figure 4: Comparison of generation quality on the LLaMA-1-13B models with BinaryMoS and OneBit.
A.5 Limitations
While BinaryMoS effectively enhances the linguistic capabilities of binarized LLMs, its application to extremely large models such as LLaMA-2-70B poses challenging due to the high training cost inherent to QAT-based strategies. Hence, to facilitate the use of BinaryMoS in such large-scale models, it may be necessary to adapt it into a PTQ-based approach or to integrate it with parameter-efficient training strategies in future work. Additionally, it is important to note that, despite the advancements brought about by by BinaryMoS, the reduction in linguistic performance of binarized LLMs coampred to their Float16 counterparts remains substantial. Consequently, to make LLM binarization practical for real-world applications, further advancements in binarization techniques are required.